OpenStudy (anonymous):

Find the exact curve where the following curve is the steepest: y= 50/ (1+6e^-2t) for t>or =to 0

OpenStudy (anonymous):

exact point or exact curve?

OpenStudy (anonymous):

sorry exact point. thanks

OpenStudy (amistre64):

wouldnt the steepest be the max slope?

OpenStudy (anonymous):

solve this equation to get the value of x and you can find y by substituting value of x to your equation: dy/dx = 0

OpenStudy (anonymous):

Well I wld say the min slope

OpenStudy (amistre64):

to find the max of a 1st derivative, derive again right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

What saruz i didn't get what u ssaid

OpenStudy (amistre64):

max or min, i was thinking physically lol

OpenStudy (anonymous):

you need to equate first derivative to zero.......you will get value of x , i.e 't' in above equaiton.

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

The first derivative seems like it never equals zero

OpenStudy (amistre64):

dy/dx would be the exaft opposite of "steepest"

OpenStudy (anonymous):

you need to solve 1st derivative=0

OpenStudy (anonymous):

it seems to be undefined at x=0 does that mke sense?

OpenStudy (anonymous):

exact opposite means @amistre64?

OpenStudy (anonymous):

I think he means that when the original function is at ii max or min then the derivtive is equal to 0?

OpenStudy (anonymous):

wait, let me solve whole equation and check!

OpenStudy (anonymous):

Thnaks I appreciate that

OpenStudy (amistre64):

right, when dy/dx = 0 we are on a completely horizontal slope

OpenStudy (amistre64):

you want the steepest slope possible, not the flatest

OpenStudy (anonymous):

ya so u find what t equals when the second derivative equals zero

OpenStudy (amistre64):

correct, at least in my understanding

OpenStudy (amistre64):

which makes sense since inflection points change slope directions

OpenStudy (anonymous):

But for some reason when I start doing it it just doesn't work out

OpenStudy (anonymous):

i gave up, how can we find steepest and not the flatest @amistre64? && @amistre64 are you from georgia tech institute?

OpenStudy (anonymous):

Well the first derivative is the slope.

OpenStudy (anonymous):

So we have to find the maximum of the slope since that is where the original function is the steepest

OpenStudy (anonymous):

So the second derivative when it equals zero that is where the first deriviative is at its max or min

OpenStudy (amistre64):

\[y= 50 (1+6e^{-2t})^{-1}\] \[y'= -50 (1+6e^{-2t})^{-2}*-12e^{-2t}\] \[y'= 600 (1+6e^{-2t})^{-2}*e^{-2t}\] \[y''= \frac{-1200 }{e^{2t}(1+6e^{-2t})^{2}}+\frac{14400}{e^{2t}\ e^{2t}(1+6e^{-2t})^{3}}\] im gonna have to chk with the wolf to see if i did the mathing right, but when tis =0 or is undefined are good places to check

OpenStudy (anonymous):

K thanks

OpenStudy (anonymous):

wolf shows no solution!!

OpenStudy (anonymous):

You may have not used enough brackets

OpenStudy (amistre64):

\[ y''=-\frac{1200 e^{2t} (e^{2t}-6))}{(e^{2t}+6)^3}\] is what the wolf gives up, that might be a simplification of mine; or something completely different ;)

OpenStudy (anonymous):

k i will play around

OpenStudy (anonymous):

But when you make it equal to zero it becomes undefined

OpenStudy (amistre64):

undefined is just as good as zero, it just means that you need to test that point

OpenStudy (anonymous):

HUH?

OpenStudy (anonymous):

So then it is undefined at (0,0)?

OpenStudy (anonymous):

i m guessing that it never touches x-axis

OpenStudy (amistre64):

derivatives are finicky. they can give false readings if we assume that dy/dx = 0 is a min or a max. all the critical points to check are: =0, or undefined

OpenStudy (anonymous):

But when it is undefined You can't solve it

OpenStudy (amistre64):

if the original equation is undefined at the point given by the derivative; then yeah ....

OpenStudy (anonymous):

See cus you gotta do ln to both sides (That doesn't make sense whtvr) and ln 0 is undefined

OpenStudy (anonymous):

you can overcome that problem by finding second derivative. if second derivative is less than zero , its max and if more than zero , its min.......(doesnt work for second degree eqn)

OpenStudy (amistre64):

im not following your reasoning here ... let me reread this

OpenStudy (anonymous):

In order to find what t is equal you must use the natural logarithm for both sides

OpenStudy (anonymous):

since t is an exponent

OpenStudy (anonymous):

and when you do ln0 it is undefined

OpenStudy (amistre64):

you mean when this = 0 right? \[y''=-\frac{1200 e^{2t} (e^{2t}-6))}{(e^{2t}+6)^3}\] which can only be zero when e^(2t) = 6

OpenStudy (amistre64):

1200e^(2t) is never zero; then denominator is never zero; so that leaves: e^(2t) - 6 = 0

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

ya so I messed up

OpenStudy (anonymous):

Thanks for ur help ;D

OpenStudy (amistre64):

;) youre welcome

OpenStudy (anonymous):

can you come help me with another question i posted

OpenStudy (anonymous):

A plot is attached.

OpenStudy (anonymous):

Thanks I didn't need help with this question but with another one I posted but Thanks neways